OpenStudy (anonymous):

Evaluate the function f(x) = 2x + 2, when x = –1. 4 ♥ –4 3 0 Check my answer Please. "♥" marks what i chose

OpenStudy (mathstudent55):

What is 2 * (-1) = ?

OpenStudy (anonymous):

Replace x with -1 2(-1)+2

OpenStudy (mathstudent55):

f(x) = 2x + 2 f(-1) = 2 * (-1) + 2 Calculate the right side above.

OpenStudy (anonymous):

-2 - 2 = 0 ?

OpenStudy (mathstudent55):

No. -2 - 2 = -4 Here the problem is f(-1) = 2 * (-1) + 2 First you do 2 * (-1) which is -2 Then you add -2 to 2 f(-1) = 2 * (-1) + 2 = -2 + 2 = 0 That means f(-1) = 0 The correct answer is choice D.